OpenStudy (anonymous):

is this equation 7x+6-4x=12+3x-8 always,sometimes, or never true. 17) combine like terms on each side of the equation. +6=3x+ 18)now solve the equation 19)the equation is always,sometimes,or never true.

OpenStudy (anonymous):

@calculusxy

OpenStudy (calculusxy):

Well I don't have a pen and paper with me now, but combine the like terms on each of the expressions between both sides. If they equal, then it will be always.

OpenStudy (anonymous):

i need help combining like terms

OpenStudy (idealist10):

7x+6-4x=12+3x-8 7x-4x+6=12-8+3x 3x+6=4+3x 3x's cancel out, 6=4 which isn't valid. So this equation is never true.

OpenStudy (midhun.madhu1987):

Just move all terms containing x to left side and u will get it..

OpenStudy (anonymous):

this is confusing

OpenStudy (midhun.madhu1987):

 Hope its clear now..